PDS_VERSION = PDS3 RECORD_TYPE = STREAM OBJECT = TEXT PUBLICATION_DATE = 1997-03-05 NOTE = "MGS RST Instrument Health Report HEA7064A.TXT" END_OBJECT = TEXT END Summary ======= Slow data transfer; SOPC sendmail running amok? Antenna motion still visible in closed loop data Data Transfer ============ Data transfer on our NASCOM line dropped this week from near capacity (56 kbs) to about a third of that. The channelized engineering data summarized above totaled almost 138 MB; it took 16 hrs 32 minutes for the transfer, an average rate of a little more than 18.5 kbs. A short test using cdb_wotu this morning yielded a similar transfer rate. A check of /var/log/syslog showed an unusual number of sendmail error messages, apparently from daemon-originated processes trying to contact unreachable hosts at JPL. We rebooted, and a quick check of data rate with cdb_wotu indicates that we have recovered at least most of normal capacity. The slow transfer was probably partly to blame for our falling behind after last weekend. Engineering Data ================ Engineering Data Statistical Summary from CSV2STAT: First SCET: 1997-03-04T00:00:04.674 Last SCET: 1997-03-04T23:46:00.170 CHANNL CHANNEL TITLE DN-LO DN-HI EU-LO EU-HI EU-AVG/STDEV RECDS ------ ------------- ----- ----- --------- --------- ---------/-------- ----- L-0037 MOT1_USO_ENA 0 0 29924 L-0047 MOT2_USO_ENA 1 1 29924 L-0053 RF_SW_OSC 2 2 29923 L-0060 TWTA1_FILMNT 0 0 29924 L-0061 TWTA1_HV 0 0 29934 L-0070 TWTA2_FILMNT 1 1 29934 L-0071 TWTA2_HV 0 0 29924 L-0080 USO_PWR 0 0 29924 L-0090 KaBLE_STAT 0 1 29924 L-0091 KaBLE_ENABLE 0 0 29934 L-0111 MOT1_RCV_AGC 253 255 -1.52E+02 -1.51E+02 -1.52E+02/6.61E-05 15067 L-0121 MOT2_RCV_AGC 116 255 -1.53E+02 -1.07E+02 -1.12E+02/7.33E-02 15067 L-0131 TWTA1_HLX_I 0 1 -3.84E-02 -1.85E-02 -3.83E-02/8.92E-06 15067 L-0132 TWTA1_ANOD_V 0 1 -1.00E+03 -1.00E+03 -1.00E+03/2.50E-03 14883 L-0141 TWTA2_HLX_I 38 43 7.05E-01 8.05E-01 7.49E-01/1.29E-04 15067 L-0142 TWTA2_ANOD_V 58 64 -8.72E+02 -8.60E+02 -8.66E+02/1.31E-02 14883 L-0200 USO_REG_V 203 205 4.06E+00 4.10E+00 4.08E+00/6.47E-05 14883 L-0201 USO_OVEN_V 123 126 2.46E+00 2.52E+00 2.48E+00/6.09E-05 14883 T-0302 EPC1_T 118 138 1.98E+01 2.85E+01 2.36E+01/3.74E-02 3739 T-0303 EPC2_T 111 142 1.81E+01 3.16E+01 2.30E+01/5.98E-02 3739 T-0310 RF_ISOLTR1_T 182 209 1.56E+01 2.75E+01 2.02E+01/2.52E-02 15066 T-0311 RF_ISOLTR2_T 170 192 2.38E+01 3.17E+01 2.72E+01/1.63E-02 15066 T-0312 TWT1_T 180 210 1.50E+01 2.82E+01 1.98E+01/5.44E-02 3739 T-0313 TWT2_T 165 189 2.49E+01 3.35E+01 2.85E+01/3.42E-02 3739 T-0315 USO_T 177 180 9.93E+00 1.09E+01 1.03E+01/2.86E-03 3739 T-0316 KA_AMP_T 151 201 1.98E+01 3.86E+01 2.47E+01/5.69E-02 7454 L-0200 remains constant at 4.08 volts. L-0201 has been steady since yesterday at 2.48 volts. T-0315 is up slightly (+0.3C) to 10.3C. Open Loop Data ============== Data Collection --------------- No new data Data Anomalies -------------- USO Test #8 data have been given quick-look processing. Spurs at +/-650 Hz from the carrier and at 60 and 180 Hz from DC remain after KaBLE was turned off. Also, the 5 minute amplitude modulation continued without any change in apparent characteristics. Analyses -------- Joe Twicken has completed analysis of data from USO Test #8 while KaBLE was on. He will submit a consolidated report when processing of the KaBLE off data has been completed. Closed Loop Data ================ Data Collection --------------- We have obtained the following files from OSCAR and EPIC: ATDF: 97049T049.SC94;1 97051T052.SC94;1 ODF: 97063C064.DAT;1 IONCAL: TROPCAL: WEATHER: WEATHER_10_LATEST.TXT;18 WEATHER_40_LATEST.TXT;18 WEATHER_60_LATEST.TXT;18 UTPM: EOP: 970304-eop.long 970304-eop.short Analyses -------- Bill Sjogren has examined tracking data for signs of antenna motion since pointing was adjusted on 1997-03-04T20:35. He has provided plots of 10-minute Doppler residuals which suggest that the HGA motion has actually increased since the correction. Results are summarized in the table below. Start Date/ERT End Date/ERT DSS Doppler Residual Oscillation Period Peak-to-Peak Amplitude ------------------- ------------------- --- -------- ------------------------ 1997-03-03T22:26:10 1997-03-04T04:07:10 65 100 min 0.006 mm/sec 1997-03-04T17:46:59 1997-03-05T00:00:59 65 100 min 0.011 mm/sec 1997-03-05T00:47:58 1997-03-05T07:12:58 15 100 min 0.013 mm/sec